an dhe icale Mncther the ieals.andhe Liobilides Ni 个 Net TIrtare

How do you determine the amount of overstated or understated categories?
Answer #1

Downstream transportation is a period expense. But the CFO has misclassified that as product cost.

Transporatation cost assigned per unit = $ 625,600 / 9,200 = $ 68 per unit produced..

But all units produced did not get sold.

Ending inventory = 9,200 - 7,800 = 1,400 units.

Therefore, the CFO managed to defer 1,400 x $ 68 = $ 95,200 of the total transportation expense to the next accounting period.

Asset ( Inventory) Overstated $ 95,200
Liabilities N/A -
Retained Earnings Overstated $ 95,200
Reserve N/A -
Expense Understated $ 95,200
Net Income Overstated $ 95,200